Assume that we have three assets. 

The first one has expected return μ1 = 10% and standard deviation of return equal to σ1 = 0.14. The second has expected return μ2 = 20% and standard deviation of return equal to σ2 = 0.2. The third asset has expected return μ3 = 15%.


We would like to determine the range of the standard deviation of the third asset so that non of the asset dominates another. 


This range is an interval with a lower bound a and an upper bound b.


What equals the lower bound a of the interval? Please insert your result with two decimals.


1
Expert's answer
2021-05-11T14:22:48-0400

Firstly, we can calculate the return per unit of risk for first asset as follows:

"Return\space per\space unit \space of\space risk=\frac{Return}{ Standard\space deviation}"

"=\frac{10}{0.14}=71.4285714285"


Now, we can calculate return per unit of risk for second asset as follows:

"=\frac{20}{0.2}=100"


Now, we can see that return per unit of risk is more for second asset as compared to first asset. In such a case second asset will dominate first asset. 

So, lower bound of standard deviation for third asset can be calculated as follows:

Risk per unit of third asset should be less than 100

"\frac{15}{Standard\space deviation} <100"


"Standard\space deviation>\frac{15}{100}"


"Standard\space deviation>0.15"


So, lower bound will be 0.15 for standard deviation of third asset because in this case, third asset and second asset will not be able to dominate each other.